# Actinide signatures in low electron fraction kilonova ejecta

### v1

@@ Abstract
pubyear: 2024pagerange: Actinide signatures in low electron fraction kilonova ejectaB
@@ 1 Introduction
The parametrization of power and composition by $Y_{e}$ can be extended to calculations of expansion opacities, which are then used to constrain groups of elements present in KN ejecta following the evolution of broadband LCs (e.g. Kasen et al., 2017; Tanvir et al., 2017; Villar et al., 2017; Tanaka et al., 2020; Domoto et al., 2021, 2022). The identification of individual elements and species requires detailed spectral analysis (e.g. Watson et al., 2019; Domoto et al., 2021; Gillanders et al., 2021; Domoto et al., 2022; Gillanders et al., 2022; Sneppen & Watson, 2023; Tarumi et al., 2023). In both cases, models making use of atomic data (e.g. at least energy levels and radiative transitions), as well as ejecta compositions and power, typically from a general formula or taken from nuclear network outputs, are required in order to interpret observational data.
The case of low $Y_{e}$, actinide bearing ejecta is particularly challenging due to the large degree of uncertainty associated with such neutron rich conditions. From the nuclear perspective, many studies have shown that synthesised abundances of actinide species are poorly constrained due to unknown nuclear quantities such as nuclear masses, $\beta$-decay rates, decay channel branching ratios, and nuclear fission rates (e.g. Mendoza-Temis et al., 2015; Wu et al., 2019; Giuliani et al., 2020; Lemaître et al., 2021; Zhu et al., 2021; Wu & Banerjee, 2022). This leads to large variations in predicted abundances, as well as radioactive power, both of which are key ingredients in the modelling of KNe. Furthermore, the atomic structure of actinide species is typically poorly known, due to a lack of experimental atomic data, and the complexity of actinide species with partially filled, open f-shell configurations (e.g. Even et al., 2020; Tanaka et al., 2020; Flörs et al., 2023; Fontes et al., 2023).
@@
2 Models
@@
Table 1: Model compositions by mass fractions. B16B and W14 are at 25 days, while B16A is at 27 days.
@@
From Figs. 1 and 2, the different outputs of the three models for an identical electron fraction are apparent. Considering first the B16B and B16A models which only differ by ground-state nuclear masses, we find significant variations in actinide production. Notably, the B16B model predicts peak actinide abundance for 94Pu, as well as significant abundances of 96Cm and 98Cf, whereas the B16A values are markedly lower (red and black curves in Fig. 1). Unfortunately, atomic data for such heavy species are lacking, such that their spectral signatures cannot be modelled directly in the radiative transfer simulation. However, these species are important starting points for energetic spontaneous fission (SF) and $\alpha$-decay channels that may dominate power production should they be present in sufficient amounts (e.g. Barnes et al., 2016; Rosswog et al., 2017; Zhu et al., 2018; Wanajo, 2018; Wu et al., 2019; Giuliani et al., 2020; Holmbeck et al., 2023).
Indeed, looking at Fig. 2, we see that the B16A model has an energy deposition that is dominated by $\alpha$-decay across the entire 1 – 100 day timespan. There are several key differences between $\alpha$ and $\beta$-decay that must be taken into account in order to understand the impact on energy deposition. For similar lifetimes, $\alpha$-decays are typically stronger energy sources than $\beta$-decays, as they usually involve chains of several $\alpha$-decays (e.g. 4 for 225Ac) that produce $\alpha$-particles with kinetic energies significantly higher than $\beta$-decay electrons, $\sim 6$ MeV cf. $\sim 0.5$ MeV. From this, the raw radioactive power of $\alpha$-decay is greater than that of $\beta$-decay. The energy deposition is then further augmented by the efficient thermalization of $\alpha$-particles arising from their greater energy and higher charges (e.g. Barnes et al., 2016; Kasen & Barnes, 2019; Wu et al., 2019).
@@
In comparison to the B16A/B models, the W14 model produces a markedly different abundance pattern, particularly for the actinides. A very low actinide mass-fraction of $X_{\rm{Ac}}\sim 10^{-3}$ is yielded, approximately an order of magnitude less than the B16A and B16B calculations. Conversely, a higher lanthanide mass fraction is obtained ($X_{\rm{La}}\sim 0.26$ cf. $\sim 0.13$ for B16A), as well as significantly higher abundances of elements ranging from 36Kr to 50Sn, thus covering parts of the first and second r-process peaks.
@@
The above discussion highlights the importance of using time-dependent, radioactive energy deposition rates with consideration of the distribution of the various decay products as determined from network outputs, in order to fully capture the dependence on nuclear properties (e.g. Barnes et al., 2021). These effects are difficult to capture by analytical heating formulae despite the inclusion of multiple terms (e.g. Rosswog & Korobkin, 2022). In addition, the thermalization of decay products plays a fundamental role in the energy deposition, a subject which has been greatly studied (Barnes et al., 2016; Wollaeger et al., 2018; Kasen & Barnes, 2019; Waxman et al., 2019; Hotokezaka & Nakar, 2020; Zhu et al., 2021). In this work, the analytical thermalization equations from Kasen & Barnes (2019) are used for $\alpha$-decay particles, $\beta$-decay electrons/positrons, and $\gamma$-rays, with an additional input to the $\alpha$ and $\beta$-decay treatments from Waxman et al. (2019), while the prescription from Barnes et al. (2016) is employed for fission products, regardless of initial injection energy (see Appendix A of Pognan et al., 2023, for a detailed description of the employed formulae). While this treatment may not be the most advanced, the focus of this work is not the uncertainty present within thermalization physics, but rather the impact of uncertainties in the nuclear network inputs, i.e. the raw radioactive power, on potential actinide signatures in KNe. As such, an equal application of a relatively simple thermalization prescription is deemed to be satisfactory for the goals of this study.
@@ 2.1 Radiative transfer simulation
[…]
@@ 3 Thermodynamic Evolution
where $\Lambda^{i}$ is the line cooling function for a given ion species, dependent on temperature $T$, ion density $n_{i}$ and electron density $n_{e}$, in complex and non-linear ways originating from the atomic structure of the species. Taking the line cooling function to be approximately independent of electron and ion density in low enough density regimes, such as those studied here, we find a cooling per volume of $c_{\rm{line}}(t)\propto t^{-6}$.
@@ 4.1 Actinide-rich and actinide-free spectra
It is interesting to find that even with relatively small actinide mass fractions of $X_{\rm{Ac}}\sim 0.018-0.030$ in the two B16A/B models, we recover significant actinide emission in the overall spectrum. As open f-shell species, actinides have a large energy level count and correspondingly numerous transitions. This allows them to efficiently absorb blue optical photons and produce redder photons through fluorescence cascades, even more effectively than the lanthanides (Flörs et al., 2023; Pognan et al., 2023). Several low lying transitions from actinide ions that are no longer open f-shell species, such as Ac iii and U iv described in the next subsection, also produce significant flux by direct emission, as opposed to fluorescence. These factors combined likely lead to the actinides providing anywhere from $\sim$ 40 – 80 per cent of the emergent model fluxes, despite their small mass fractions. There is naturally a limit however, and we see that the very low actinide mass fraction of $X_{\rm{Ac}}\sim 10^{-3}$ in the W14 model is insufficient to provide significant actinide emission in the emergent spectra.
@@
4.2 Actinide contributions to the emergent spectra
The contribution of individual actinide species, as well as the optically thick lines present in the models, at selected epochs can be viewed in Figs. 6 to 9. Since the W14 model does not exhibit marked actinide features, we focus here on the B16A and B16B models, while the spectral evolution of the W14 model can be viewed in Appendix B.
Considering first Fig. 6, we have the B16A model spectra in the left-hand panels, and the corresponding optically thick lines in the right-hand panels, from 11 to 27 days. The optically thick line plots show optically thick ($\tau\geq 1$) transitions in the 5 model zones from velocity coordinates $v_{\rm{inner}}=0.05$c to $v_{\rm{outer}}=0.3$c, in steps of 0.05c. The height of the markers in the plot shows to which point in the ejecta the transition is optically thick e.g. if the marker reaches 0.2c, the transition is optically thick out until the third ejecta layer. In the analysis of the spectra below, we use the optically thick line plots to understand where absorption troughs in the spectra originate from.
Looking at the 11 days epoch in the top panels, the most active actinide species in spectral formation are found to be doubly and triply ionized 91Pa and 92U, with some weaker contributions from 90Th. The main actinide-dominated features here are the 92U P-Cygni like features at $\sim 4000$ and $\sim 8000\textup{\AA}$ respectively, and the blended 91Pa and 92U feature at 1 micron. The double feature from Pa iii and Pa iv at 1.7 micron contributes significantly to the emergent spectral shape at this wavelength, but does not entirely dominate, being heavily blended with lanthanide contributions (see Fig. 4). The U iv transitions, as well as the Pa iv transition at 1 micron, are found to be strong ($A\gtrsim 10^{5}\,\rm{s^{-1}}$) allowed transitions to the ground state. Conversely, the blended 91Pa feature at 1.7 micron results from the contribution of many diverse allowed transitions with smaller radiative transition rates of $A\lesssim 10^{4}\,\rm{s^{-1}}$, typically seen in fluorescence cascades.
@@ 4.2 Actinide contributions to the emergent spectra
The U iv emission at 4.4 micron in the B16B model, which dominates flux contribution there past $\sim 30$ days is somewhat reminiscent of the 4.5 micron Spitzer detection at 43 and 74 days for AT2017gfo (Kasliwal et al., 2022), which has been tentatively attributed to either Se iii or W iii by also taking into account the lack of detection at 3.6 micron (Hotokezaka et al., 2022). Looking at Table 1, we see that our models include 74W, and in significant abundances for the B16A and W14 models. The relevant transitions of W iii for 4.5 micron emission are the $n=3\rightarrow 2$ and $n=2\rightarrow 1$ forbidden transitions with $\lambda=4.535\mu$m and $\lambda=4.432\mu$m respectively, following data available from the National Institute of Standards and Technology (NIST) Atomic Spectral Database (Kramida et al., 2020). The fac data for W iii have not been calibrated for this study however, and have these transitions at $\lambda=4.75\mu$m and $5.22\mu$m respectively. The former lies within the spectral range, and therefore may be expected to produce noticeable emission in the emergent spectra. Looking at the W14 model spectra in Appendix B, we see some W iii emission at $\sim 4.5$ micron appearing, though it remains largely subdominant compared to lanthanide emission at the same wavelength. In the B16A/B models however, no significant W iii emission is obtained.
@@
This may be so for a variety of reasons. For forbidden transitions to be prominent, we require a significant abundance of the species, i.e. by elemental abundance and ion fraction, as well as the collisional de-excitation time-scale to be longer than the radiative transition lifetime, which is found to be on the order of $\sim 5$ s in our data. Considering first the B16A model at 41 days, we find that W iii is present mostly in the innermost ejecta layer with, but a relatively low ion fraction of $\sim 0.07$. Conversely, the B16B model at 47 days has a larger W iii ion fraction of $\sim 0.27$, but a much lower 74W abundance in the model composition to begin with (see Table 1). As such, the W iii emission in both these models may be quite low in comparison to other species. It is also possible that the conditions for W iii forbidden emission are simply not met in the B16A/B models, requiring a balance between low density for inefficient collisions, but also a degree of ionization that favours W iii over W iv. Models with lower energy depositions than those yielded by dominant $\alpha$-decay, and therefore yielding significant ion fractions of W iii in low density ejecta, may be required in order for strong emission from this species to be produced.
@@
The Th iii emission in our models may be checked by comparing our energy level structure to data available from NIST. In doing so, we find that the wavelengths of our prominent Th iii transitions are quite inaccurate, e.g. the emission at $\lambda_{0}=3.55\mu$m is in reality found to be at $13.0\mu$m, while the $\lambda_{0}=3.01\mu$m emission is at $22.3\mu$m. By also considering the change in radiative transition rate (A-value) arising from shifted energy levels, $A\propto\left(\Delta E_{\rm{NIST}}/\Delta E_{\rm{FAC}}\right)^{3}$ for dipole transitions (see e.g. Mulholland et al., 2024), we find A-values for the transitions at their correct wavelengths of $A_{13.0\mu m}\sim 1.7\times 10^{2}\,\rm{s^{-1}}$ and $A_{22.3\mu m}\sim 2.9\times 10^{1}\,\rm{s^{-1}}$ respectively. While this does mean that the Th iii emission in our model spectra cannot be taken directly as potentially observable actinide signatures, these results suggest that Th iii may produce significant mid-IR emission, should the temperature and ionization structure of the ejecta be comparable to those of the B16B model found here. The $22.3~{}\mu$m transition is particularly interesting, as it corresponds to the $n=3\rightarrow 2$ transition from the lowest lying $\rm{5f^{1}6d^{1}}$ state to the lowest lying $\rm{6d^{2}}$ state, and may therefore be a good candidate for late-time, nebular phase observations with JWST, given the relatively high A-value and low excitation energies of the involved levels.
At later times $\gtrsim 60$ days, the B16A model develops two Ac iii features at $\sim 3200,4400\textup{\AA}$ respectively, which blend with the U iii and U iv features, as well as emission from Tm iii and Eu iii, and Sm iii respectively. However, the $4400~{}\textup{\AA}$ emission from Ac iii provides approximately half of the actinide flux within the $\sim 5000\textup{\AA}$ emission feature. These Ac iii emission features are worthy of note, as the atomic data underlying them have wavelengths accurate to within 2 per cent, and A-values within a factor of 2, of data available from NIST. As such, these features are robust predictions that may remain important even if the other features in that area, of which the accuracy is difficult to verify, are corrected to other wavelengths or flux levels. However, it is difficult to predict whether they will become more or less dominant in that regime if the calibration of other atomic species yields a greater number of important transitions in the same wavelength range.
The verification of the features arising from U iv is more complicated, as no level data is readily available from NIST. The three features that are found to be significant in our model spectra arise from transitions between low-lying even parity $\rm{5f^{2}6d^{1}}$ states, to odd parity $\rm{5f^{3}}$ states. Several experimental measurements (Carnall & Crosswhite, 1985; Andres et al., 1996) and Ab initio theoretical calculations (Ruipérez et al., 2007; Roy & Prasad, 2020) of low-lying $\rm{5f^{3}}$ levels in U iv are available, while only a single theoretical calculation for $\rm{5f^{2}6d^{1}}$ levels was found (Seijo & Barandiarán, 2003).

### v2

@@ Abstract
pubyear: 2025pagerange: Actinide signatures in low electron fraction kilonova ejectaB
@@ 1 Introduction
The parametrization of power and composition by $Y_{e}$ can be extended to calculations of expansion opacities, which are then used to constrain groups of elements present in KN ejecta following the evolution of broadband LCs (e.g. Kasen et al., 2017; Tanvir et al., 2017; Villar et al., 2017; Banerjee et al., 2020; Tanaka et al., 2020; Domoto et al., 2021, 2022; Banerjee et al., 2024). The identification of individual elements and species requires detailed spectral analysis (e.g. Watson et al., 2019; Domoto et al., 2021; Gillanders et al., 2021; Domoto et al., 2022; Gillanders et al., 2022; Sneppen & Watson, 2023; Tarumi et al., 2023). In both cases, models making use of atomic data (e.g. at least energy levels and radiative transitions), as well as ejecta compositions and power, typically from a general formula or taken from nuclear network outputs, are required in order to interpret observational data.
The case of low $Y_{e}$, actinide bearing ejecta is particularly challenging due to the large degree of uncertainty associated with such neutron rich conditions. From the nuclear perspective, many studies have shown that synthesised abundances of actinide species are poorly constrained due to unknown nuclear quantities such as nuclear masses, $\beta$-decay rates, decay channel branching ratios, and nuclear fission rates (e.g. Mendoza-Temis et al., 2015; Wu et al., 2019; Giuliani et al., 2020; Lemaître et al., 2021; Zhu et al., 2021; Wu & Banerjee, 2022). This leads to large variations in predicted abundances, as well as radioactive power, both of which are key ingredients in the modelling of KNe. Furthermore, the atomic structure of actinide species is typically poorly known, due to a lack of experimental atomic data, and the complexity of actinide species with partially filled, open f-shell configurations (e.g. Even et al., 2020; Tanaka et al., 2020; Flörs et al., 2023; Fontes et al., 2023; Domoto et al., 2024).
@@
2 Ejecta models and radiative transfer simulation
2.1 Ejecta models and compositions
@@
Table 1: Model compositions by mass fractions, with total lanthanide and actinide mass fractions. B16B and W14 are at 25 days, while B16A is at 27 days.
@@
From Figs. 1 and 2, the different outputs of the three models for an identical electron fraction are apparent. Considering first the B16B and B16A models which only differ by ground-state nuclear masses, we find significant variations in actinide production. Notably, the B16B model predicts peak actinide abundance for 94Pu, as well as significant abundances of 96Cm and 98Cf, whereas the B16A values are markedly lower (red and black curves in Fig. 1). Unfortunately, atomic data for such heavy species are lacking, such that their spectral signatures cannot be modelled directly in the radiative transfer simulation. However, these species are important starting points for energetic spontaneous fission (SF) and $\alpha$-decay channels that may dominate power production should they be present in sufficient amounts (e.g. Barnes et al., 2016; Rosswog et al., 2017; Zhu et al., 2018; Wanajo, 2018; Wu et al., 2019; Giuliani et al., 2020; Holmbeck et al., 2023). The total actinide mass fraction is also significantly higher in the B16B model at $X_{\rm{Ac}}=0.0306$ compared to the B16A model at $X_{\rm{Ac}}=0.0183$.
Looking at Fig. 2, we see that the B16A model has an energy deposition that is dominated by $\alpha$-decay across the entire 1 – 100 day timespan. There are several key differences between $\alpha$ and $\beta$-decay that must be taken into account in order to understand the impact on energy deposition. For similar lifetimes, $\alpha$-decays are typically stronger energy sources than $\beta$-decays, as they usually involve chains of several $\alpha$-decays (e.g. 4 for 225Ac) that produce $\alpha$-particles with kinetic energies significantly higher than $\beta$-decay electrons, $\sim 6$ MeV cf. $\sim 0.5$ MeV. From this, the raw radioactive power of $\alpha$-decay is greater than that of $\beta$-decay. The energy deposition is then further augmented by the efficient thermalization of $\alpha$-particles arising from their greater energy and higher charges (e.g. Barnes et al., 2016; Kasen & Barnes, 2019; Wu et al., 2019).
@@
In comparison to the B16A/B models, the W14 model produces a markedly different abundance pattern, particularly for the actinides. A very low actinide mass-fraction of $X_{\rm{Ac}}\sim 10^{-3}$ is yielded, approximately an order of magnitude less than the B16A and B16B calculations. Conversely, a higher lanthanide mass fraction is obtained ($X_{\rm{La}}\sim 0.26$ cf. $\sim 0.13$ for B16A and $\sim 0.07$ for B16B), as well as significantly higher abundances of elements ranging from 36Kr to 50Sn, thus covering parts of the first and second r-process peaks.
@@
The above discussion highlights the importance of using time-dependent, radioactive energy deposition rates with consideration of the distribution of the various decay products as determined from network outputs, in order to fully capture the dependence on nuclear properties (e.g. Barnes et al., 2021). These effects are difficult to capture by analytical heating formulae despite the inclusion of multiple terms (e.g. Rosswog & Korobkin, 2024). In addition, the thermalization of decay products plays a fundamental role in the energy deposition, a subject which has been greatly studied (Barnes et al., 2016; Wollaeger et al., 2018; Kasen & Barnes, 2019; Waxman et al., 2019; Hotokezaka & Nakar, 2020; Zhu et al., 2021). In this work, the analytical thermalization equations from Kasen & Barnes (2019) are used for $\alpha$-decay particles, $\beta$-decay electrons/positrons, and $\gamma$-rays, with an additional input to the $\alpha$ and $\beta$-decay treatments from Waxman et al. (2019), while the prescription from Barnes et al. (2016) is employed for fission products, regardless of initial injection energy (see Appendix A of Pognan et al., 2023, for a detailed description of the employed formulae). While this treatment may not be the most advanced, the focus of this work is not the uncertainty present within thermalization physics, but rather the impact of uncertainties in the nuclear network inputs, i.e. the raw radioactive power, on potential actinide signatures in KNe. As such, an equal application of a relatively simple thermalization prescription is deemed to be satisfactory for the goals of this study.
@@ 2.2 Radiative transfer simulation
The atomic data used in this study comes from calculations done using the Flexible Atomic Code (fac Gu, 2008, version from April 21st, 2021111Corresponding to commit id ac5f58c.), and are identical to those used in Pognan et al. (2023), the purpose of which was completeness rather than detailed accuracy for key species. The configurations used for the fac calculations can be found in the appendix of Pognan et al. (2023)  [^2] See erratum of Pognan et al. (2023) concerning the configurations of U iv., though it should be noted that the development of the methods implemented in fac has caused variations in the resultant atomic data over the years from seemingly similar models. This should be considered when making detailed comparisons between data sets computed with different versions.
@@ 3 Thermodynamic Evolution
where $\Lambda^{i}$ is the line cooling function for a given ion species, dependent on temperature $T$, ion density $n_{i}$ and electron density $n_{e}$, in complex and non-linear ways originating from the atomic structure of the species. As the ion and electron densities scale as $t^{-3}$ in spherical homologous expansion ($V\propto v_{\rm{ej}}^{3}t^{3}$), and taking the line cooling function to be approximately independent of these densities in low enough density regimes, such as those studied here, we find a cooling per volume of $c_{\rm{line}}(t)\propto t^{-6}$.
@@ 4.1 Actinide-rich and actinide-free spectra
It is interesting to note that for the relatively late epochs studied here of $\sim$10 – 100 days, an actinide mass fraction of $10^{-3}$ is insufficient to produce notable spectral signatures, while a much smaller mass fraction of $10^{-5}$ was found to still produce significant absorption from Th iii at very early (1.5 – 2.5 days) times (Domoto et al., 2024). In that case however, the lanthanide fraction was also considerably smaller than in the W14 model, at $X_{\rm{La}}\sim 6\times 10^{-4}$ cf. $X_{\rm{La}}\sim 2\times 10^{-2}$. This implies that actinides may still produce significant spectral features even at very low abundances, provided the lanthanide abundance is not vastly superior. These lanthanides may suppress actinide features when sufficiently abundant, as is likely the case in our W14 model. The notably different ejecta conditions between early photospheric epochs, and the later nebular phase studied here, should however be kept in mind when making direct comparisons, as the ejecta conditions are highly different and therefore may potentially lead to equally different emergent spectra.
Following the above argument, it is perhaps not entirely surprising to find that even with relatively small actinide mass fractions of $X_{\rm{Ac}}\sim 0.018-0.030$ in the two B16A/B models, we recover significant actinide emission in the overall spectrum. As open f-shell species, actinides have a large energy level count and correspondingly numerous transitions. This allows them to efficiently absorb blue optical photons and produce redder photons through fluorescence cascades, even more effectively than the lanthanides (Flörs et al., 2023; Fontes et al., 2023; Pognan et al., 2023). Several low lying transitions from actinide ions that are no longer open f-shell species, such as Ac iii and U iv described in the next subsection, also produce significant flux by direct emission, as opposed to fluorescence. These factors combined likely lead to the actinides providing anywhere from $\sim$ 40 – 80 per cent of the emergent model fluxes, despite their small mass fractions.
@@
4.2 Spectral features and evolution of the actinide-rich models
The contribution of individual actinide species, as well as the optically thick lines present in the models, at selected epochs can be viewed in Figs. 6 to 9. n these figures, the model spectra are shown in the left-hand panels, and the corresponding optically thick lines in the right-hand panels. The line plots show optically thick ($\tau\geq 1$) transitions in the 5 model zones from velocity coordinates $v_{\rm{inner}}=0.05$c to $v_{\rm{outer}}=0.3$c, in steps of 0.05c. The height of the markers in the plot shows to which point in the ejecta the transition is optically thick e.g. if the marker reaches 0.2c, the transition is optically thick out until the third ejecta layer. In the analysis of the spectra below, we use the optically thick line plots to understand where absorption troughs in the spectra originate from. Since the W14 model does not exhibit marked actinide features, we focus here on the B16A and B16B models, while the spectral evolution of the W14 model can be viewed in Appendix B.
4.2.1 Early times: 10 - 30 days
Considering first Fig. 6, we have the B16A model spectra and optically thick lines from 11 to 27d. Looking at the 11 days epoch in the top panels, the most active actinide species in spectral formation are found to be doubly and triply ionized 91Pa and 92U, with some weaker contributions from 90Th. The main actinide-dominated features here are the 92U P-Cygni like features at $\sim 4000$ and $\sim 8000\textup{\AA}$ respectively, and the blended 91Pa and 92U feature at 1 micron. The double feature from Pa iii and Pa iv at 1.7 micron contributes significantly to the emergent spectral shape at this wavelength, but does not entirely dominate, being heavily blended with lanthanide contributions (see Fig. 4). The U iv transitions, as well as the Pa iv transition at 1 micron, are found to be strong ($A\gtrsim 10^{5}\,\rm{s^{-1}}$) allowed transitions to the ground state. Conversely, the blended 91Pa feature at 1.7 micron results from the contribution of many diverse allowed transitions with smaller radiative transition rates of $A\lesssim 10^{4}\,\rm{s^{-1}}$, typically seen in fluorescence cascades.
@@ 4.2.1 Early times: 10 - 30 days
4.2.2 Late times: 40 - 100 days
Table 2: Emitted flux of W iii at 4.75 and 5.22 micron compared to the total emitted flux of the models, in units of $\rm{erg\,s^{-1}\,cm^{-2}\,\textup{\AA}^{-1}}$. The B16A model is at 41 days, and the B16B/W14 models at 47 days.
B16A B16B W14
$\lambda$ [$\mu$m] $F_{\lambda}^{\rm{tot}}$ $F_{\lambda}^{\rm{W\,\textsc{iii}}}$ $F_{\lambda}^{\rm{W\,\textsc{iii}}}/F_{\lambda}^{\rm{tot}}$ $F_{\lambda}^{\rm{tot}}$ $F_{\lambda}^{\rm{W\,\textsc{iii}}}$ $F_{\lambda}^{\rm{W\,\textsc{iii}}}/F_{\lambda}^{\rm{tot}}$ $F_{\lambda}^{\rm{tot}}$ $F_{\lambda}^{\rm{W\,\textsc{iii}}}$ $F_{\lambda}^{\rm{W\,\textsc{iii}}}/F_{\lambda}^{\rm{tot}}$
4.75 $8.448\times 10^{-20}$ $7.296\times 10^{-21}$ 0.086 $2.488\times 10^{-20}$ $4.270\times 10^{-22}$ 0.017 $3.562\times 10^{-20}$ $3.059\times 10^{-21}$ 0.086
5.22 $4.534\times 10^{-20}$ $7.912\times 10^{-21}$ 0.175 $1.723\times 10^{-20}$ $5.020\times 10^{-22}$ 0.029 $3.257\times 10^{-20}$ $3.828\times 10^{-21}$ 0.117
At later times $\gtrsim 60$ days, the B16A model develops two Ac iii features at $\sim 3200,4400\textup{\AA}$ respectively, which blend with the U iii and U iv features, as well as emission from Tm iii and Eu iii, and Sm iii respectively. However, the $4400~{}\textup{\AA}$ emission from Ac iii provides approximately half of the actinide flux within the $\sim 5000\textup{\AA}$ emission feature. These Ac iii emission features are worthy of note, as the atomic data underlying them have wavelengths accurate to within 2 per cent, and A-values within a factor of 2, of data available from NIST. As such, these features are robust predictions that may remain important even if the other features in that area, of which the accuracy is difficult to verify, are corrected to other wavelengths or flux levels. However, it is difficult to predict whether they will become more or less dominant in that regime if the calibration of other atomic species yields a greater number of important transitions in the same wavelength range, though they are expected to produce strong optical features also at early times (Domoto et al., 2024).
4.2.3 Spectrally significant actinide features
The Th iii emission in our models may be checked by comparing our energy level structure to data available from NIST. In doing so, we find that the wavelengths of our prominent Th iii transitions are quite inaccurate, e.g. the emission at $\lambda_{0}=3.55\mu$m is in reality found to be at $13.0\mu$m, while the $\lambda_{0}=3.01\mu$m emission is at $22.3\mu$m. By also considering the change in radiative transition rate (A-value) arising from shifted energy levels, $A\propto\left(\Delta E_{\rm{NIST}}/\Delta E_{\rm{FAC}}\right)^{3}$ for dipole transitions (see e.g. Mulholland et al., 2024), we find A-values for the transitions at their correct wavelengths of $A_{13.0\mu m}\sim 1.7\times 10^{2}\,\rm{s^{-1}}$ and $A_{22.3\mu m}\sim 2.9\times 10^{1}\,\rm{s^{-1}}$ respectively. While this does mean that the Th iii emission in our model spectra cannot be taken directly as potentially observable actinide signatures, these results suggest that Th iii may produce significant mid-IR emission, should the temperature and ionization structure of the ejecta be comparable to those of the B16B model found here. The $22.3~{}\mu$m transition is particularly interesting, as it corresponds to the $n=3\rightarrow 2$ transition from the lowest lying $\rm{5f^{1}6d^{1}}$ state to the lowest lying $\rm{6d^{2}}$ state, and may therefore be a good candidate for late-time, nebular phase observations with JWST, given the relatively high A-value and low excitation energies of the involved levels. It would also be interesting to verify if absorption from Th iii at 1.8 micron (Domoto et al., 2024) persists into the nebular phase, though it is probable that the lines responsible for this absorption in the photospheric phase become optically thin at later times.
@@ 4.2.3 Spectrally significant actinide features
The U iv emission at 4.4 micron in the B16B model, which dominates flux contribution there past $\sim 30$ days is somewhat reminiscent of the 4.5 micron Spitzer detection at 43 and 74 days for AT2017gfo (Kasliwal et al., 2022), which has been tentatively attributed to either Se iii or W iii by also taking into account the lack of detection at 3.6 micron (Hotokezaka et al., 2022). Looking at Table 1, we see that our models include 74W, and in significant abundances for the B16A and W14 models. The relevant transitions of W iii for 4.5 micron emission are the $n=3\rightarrow 2$ and $n=2\rightarrow 1$ forbidden transitions with $\lambda=4.54\mu$m, $A=0.61\,\rm{s^{-1}}$ and $\lambda=4.43\mu$m, $A=0.62\,\rm{s^{-1}}$ respectively, following data available from the National Institute of Standards and Technology (NIST) Atomic Spectral Database (Kramida et al., 2020) and Hotokezaka et al. (2022). The fac data for W iii have not been calibrated for this study however, and have these transitions at $\lambda=4.75\mu$m, $A=0.46\,\rm{s^{-1}}$ and $5.22\mu$m, $A=0.31\,\rm{s^{-1}}$ respectively. The former lies within the spectral range, and therefore may be expected to produce noticeable emission in the emergent spectra, which, however, does not appear to be the case.
@@
This may be so for a variety of reasons. For forbidden transitions to be prominent, we require a balance between low ejecta density for inefficient collisions, but also a degree of ionization that favours W iii over W iv, i.e. density cannot be so low that recombinations are entirely inefficient. We also need a significant abundance of the species, by elemental composition of the ejecta and ion fraction, as well as the collisional de-excitation time-scale to be longer than the radiative transition lifetime, which is found to be on the order of $\sim 5$ s in our data. Considering first the B16A model at 41 days, we find that W iii is most abundant in the innermost layer, but with a low ion fraction of only $\sim 0.07$, while the subsequent outer layers completely favour W iv. Conversely, the B16B model at 47 days has a larger W iii ion fraction of $\sim 0.27$, but a much lower 74W abundance in the model composition to begin with (see Table 1). As such, the W iii emission in both these models may be intrinsically quite low in comparison to other species.
@@
We verify this by considering the contribution of W iii to the total emission at 4.75 and 5.22 micron, shown in Table 2. This emission consists of the photons emitted before radiative transfer is conducted, and physically corresponds to the emergent spectrum in a completely optically thin medium. Therefore, transfer effects such as absorption and scattering are not taken into account in this ‘optically thin’ spectrum, but these effects are expected to be mostly negligible at IR wavelengths where the ejecta are optically thin. It should be noted that the contribution of individual transitions to the W iii emission at a given wavelength cannot be extracted. For instance, the $n=4\rightarrow 3$ ground state transition in the fac data has a wavelength of 5.29 micron, such that it contributes to the recorded W iii flux at 5.22 micron.
[…]
In contrast, U iv dominates emission in the B16B model at 4.5 micron over a broad range of epochs. The verification of the features arising from U iv is more involved, as no level data is readily available from NIST. The three features that are found to be significant in our model spectra arise from transitions between low-lying even parity $\rm{5f^{2}6d^{1}}$ states, to odd parity $\rm{5f^{3}}$ states. Several experimental measurements (Carnall & Crosswhite, 1985; Andres et al., 1996) and Ab initio theoretical calculations (Ruipérez et al., 2007; Roy & Prasad, 2020) of low-lying $\rm{5f^{3}}$ levels in U iv are available, while only a single theoretical calculation for $\rm{5f^{2}6d^{1}}$ levels was found (Seijo & Barandiarán, 2003).
@@
Figure 10: Broadband lightcurves of the models in the UV (top left panel), optical (top right panel) and IR (bottom panels). The points correspond to the B16A model, the crosses to the B16B model, and the ‘Y’-shaped markers to the W14 model.